# Prefixing Attention Sinks can Mitigate Activation Outliers for Large Language Model Quantization

### v1

@@ 2 Related Work
Notably, Xiao et al. (2023) also provides two options for per-tensor activation quantization: one with dynamic quantization range, and another with static range. While these options tend to be faster than per-token (with static range being the fastest), their generation quality is much lower than per-token, especially on recent models such as LLaMA-3 (Touvron et al., 2023).
@@ 6 Analysis
  • Did the CushionCache really replaced the role of attention sink? (Section 6.2)
@@ 6.1 Change of Activation Magnitudes
In Fig. 2, we visualize the top-1/2/3 activations and median for each layer of LLaMA3-8B. The left panel shows the magnitude of outliers that occur during the standard operation of LLaMA3-8B. The median is almost zero, indicating that most values are close to zero, with only a few significantly large outliers. On the right panel, we present the same measurements after applying the CushionCache. We conclude that CushionCache effectively removes the activation outliers.

### v2

@@ 2 Related Work
Notably, Xiao et al. (2023) also provides two options for per-tensor activation quantization: one with dynamic quantization range, and another with static range. While these options tend to be faster than per-token (with static range being the fastest), their generation quality is much lower than per-token, especially on recent models such as LLaMA3 (Touvron et al., 2023).
@@ 5.3 Ablation Study
5.5 Other experiments
In the Appendix A, we provide additional experiments results on the following topics:
@@ 6 Analysis
  • Did the CushionCache really replace the role of attention sink? (Section 6.2)
@@ 6.1 Change of Activation Magnitudes
In Fig. 2, we visualize the top-1/2/3 activations and median for each layer of LLaMA3-8B. The left panel plots the magnitude of the median and top-3 activations that occur during the standard operation of LLaMA3-8B. We observe that the median is almost zero, indicating that a significant fraction of all activations are close to zero, with only a few significantly large outliers. On the right panel, we plot the same values after applying the proposed CushionCache algorithm. We observe that the size of the top-3 activations have dramatically decreased, leading to a conclusion that CushionCache effectively removes the activation outliers.
@@ 7 Conclusion
Acknowledgements
This work was supported in part by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No. RS-2023-00213710), and in part by the Korea government (MSIT) (No. RS-2024-00453301). JL also thanks Hong-Seok Kim and Radha Chandika for their generous support during his visit at Google.
Appendix A Additional experiments
In this section, we provide additional experiments that have been missing in the main script.
A.1 MMLU dataset
We have additionally evaluated the quantized model on MMLU dataset, which encompasses a much larger set of tasks including STEM (Table 7). The results suggest that CushionCache remains to be effective on MMLU as well.
Model LLaMA2-7B Mistral-7B LLaMA3-8B
FP16 41.27 58.63 62.13
SmoothQuant-O3 23.76 23.62 25.32
+ CushionCache (ours) 38.06 (+14.30) 56.59 (+32.97) 58.99 (+33.67)
SmoothQuant-O2 27.88 25.66 30.94
+ CushionCache (ours) 40.45 (+12.59) 58.05 (+32.39) 60.59 (+29.65)
SmoothQuant-O1 40.76 58.70 61.88
+ CushionCache (ours) 41.65 (+0.89) 59.20 (+0.50) 61.55 (-0.33)
Table 7: Results on MMLU. We compare the results on the MMLU dataset.
A.2 Generation latency
We have measured the average latency of generating each token. We have experimented with W8A8-quantized LLaMA-3B, using the SmoothQuant kernel on a single A6000 GPU; not that this may not be the best hardware-optimized kernel for our hardware, but can be meaningful in terms of providing a comparison. We have used the prompts of length 500, and averaged over 1000 generated tokens. We compare both the time to the first token (TTFT; prefill phase) and the time per output token (TPOT; generation phase). in Table 8.
We observe that CushionCache only adds negligible latency, while enabling a much better adoption of the per-tensor static-range quantization techniques which provides a much faster decoding. In particular, we observe that adding CushionCache adds only 0.01–0.3ms in TTFT or TPOT, which is less than 0.5% of the total latency. Furthermore, as the CushionCache makes the faster option (e.g., per-tensor static) a viable option, it can even be viewed as enabling an overall speedup up to a few milliseconds.
LLaMA3-8B TTFT (ms) TPOT (ms)
Per-Tensor Static 78.01 48.71 $\pm$ 1.58
+ CushionCache (ours) 78.22 48.86 $\pm$ 0.55
Per-Tensor Dynamic 81.52 50.56 $\pm$ 0.71
+ CushionCache (ours) 81.56 50.93 $\pm$ 0.82
Per-Token Dynamic 83.35 51.75 $\pm$ 1.22
+ CushionCache (ours) 83.64 51.76 $\pm$ 0.95
Table 8: Generation latency. We measure the generation speed on LLaMA3-8B.
A.3 Other quantization methods
We have conducted additional experiments to combine with quantization algorithms other than SmoothQuant. In particular, we have conducted experiments on the following recent methods:
  • AWQ-4bit (Lin et al., 2024): A recent weight-only quantization algorithm, for demonstrating that CushionCache effectively boosts the performance of weight quantization algorithms.
  • QuaRot-4bit (Ashkboos et al., 2024): A recent weight+activation+cache quantization algorithm, for demonstrating that CushionCache works well with SOTA quantization algorithms.
  • KIVI-2bit (Liu et al., 2024): A recent KV cache quantization algorithm, for demonstrating that the KV cache of the CushionCache-quantized model can be compressed well with KV cache quantization methods.
The results are given in Table 9, where we observe that the CushionCache is indeed versatile, being able to be combined well with many different quantization methods. Note that for KIVI, we measure the GSM8K results, as the original paper does not report perplexity.
LLaMA3-8B WikiText-2 Perplexity
FP16 6.13
AWQ 6.18
+ CushionCache 6.15
+ Per-Cushion Static 8.40
+ Per-Cushion Static + CushionCache 7.01
QuaRot 8.21
+ CushionCache 7.41
LLaMA3-8B GSM8K (%)
FP16 48.75
+ KIVI 42.61
Per-tensor Static 0.06
+ KIVI 0.03
+ KIVI + CushionCache 38.29
Table 9: Other quantization methods. We combine CushionCache with AWQ, QuaRot, and KIVI.